# FedRef: Communication-Efficient Bayesian Fine-Tuning using a Reference Model

### v3

@@ I Introduction
In the related work section, we discuss FedAvg, FedProx, and FedOpt, which are existing approaches for model optimization in FL scenarios, along with their limitations regarding client-side computational costs and catastrophic forgetting [mccloskey1989catastrophic][french1999catastrophic] [goodfellow2013empirical]. We also explain Bayesian Transfer Learning. We then present FedRef, a novel, communication-efficient Bayesian fine-tuning approach that leverages a reference model. FedRef mitigates catastrophic forgetting in each round by integrating features from previous models while performing maximum a posteriori (MAP) estimation.
In earlier work, MAP estimation has been used to optimize an objective function for transfer learning between a target model with pre-trained parameters and a source model [bayesian]. In contrast, traditional model optimization in FL scenarios employs a proximal term to minimize the deviation of a local model from the global model [fedprox][reddi2021adaptivefederatedoptimization]. This proximal term can also be considered when performing MAP estimation. In our research, we use MAP estimation to optimize an objective function that is updated using aggregated model information from previous rounds exclusively on the server side, thereby minimizing the computational costs for clients.
@@
For the segmentation task, we present the predictive model performance across different rounds using graphical representations, based on evaluation metrics commonly used in medical image segmentation studies [karimi2019reducing, chen2021transunet, hatamizadeh2022unetr, zhu2022medical]: mean Intersection over Union (mIoU), Dice Coefficient, and Hausdorff Distance.
In summary, the proposed FedRef method demonstrates high predictive performance on both medical and general computer vision tasks while minimizing computational costs for clients.
Figure 1: Basic medical FedRef system.
II Related Work
@@ II-3 FedOpt
The authors of [reddi2021adaptivefederatedoptimization] focus on integrating adaptive optimizers such as Adam, Yogi, and Adagrad at the server side, leading to variants of FedAvg that are known as FedAdam, FedYogi, and FedAdagrad, respectively. Following the FedAdagrad equation (4):
@@
$g_{r}$: averaged update from all participating clients, $G_{r}$: accumulated sum of squared gradients (maintained on the server), $\tau$: small constant for numerical stability (e.g., $10^{-6}$). Following the FedAdam equation (II-3):
@@ II-4 Bayesian Transfer Learning
As a result, this paper investigates the optimization of transfer learning in an FL scenario. Whereas conventional transfer learning happens between only two models – one a pre-trained model intended for transfer, and the other a target mode – in FL scenarios, the server must aggregate and optimize many different client models. Therefore, we need a new MAP estimation problem formulation for FL scenarios.
@@ III FedRef: Communication-Efficient Bayesian Fine Tuning using a Reference Model
For optimal model performance and low client-side computational cost, we propose FedRef, a communication-efficient Bayesian fine-tuning method with a reference model that overcomes catastrophic forgetting by inferring from global models of previous rounds.
@@
The reference model, in Fig. 2, perform as a standard(target) model of proximal term objective reference to mitigate a deviation from previous round models. In the existing study [bayesian], transferred learning is a concept of model distillation and focus on how do distillate to transfer features from alternative reference models to a target model. The most of FL scenario also focus on how do distillate to transfer features from each clients to global model. Therefore, in the most of FL scenarios, the ideal situation is to the client’s model inference as an independent event and when the model inference of all clients is, on average, the most accurate phase.
@@
$D_{ref}$ is defined solely to explain our optimal MAP problem in Equation (III). This approach helps overcome catastrophic forgetting in each round by integrating features from previous rounds into the MAP problem and subsequently optimizing the integrated MAP value.
Finally, the objective function derived in Equation (11) can be expressed by combining Equations (8, 9, and 10) to form Equation (III).
@@
In Equation (III), the constant term $\sum_{k}^{K}F_{k}$ denotes the sum of client losses. The diagonal matrix $\mathrm{diag}(W_{1},\dots,W_{K})$ represents the aggregation weights (e.g., $\frac{n_{k}}{n}$). The $\theta^{1}$ is a global model and $\theta^{2}$ is a reference model. The term $(\theta_{r}-\theta_{r-1})^{2}$ signifies the $L_{2}$ regularization of the model parameters $\theta_{r}-\theta_{r-1}$.
@@
We describe FedRef in detail in Algorithm 5:
@@ IV-A1 Global
Within Python, the Flower framework enables the easy construction of federated learning servers and clients from high-level code. It supports both Object-Oriented Programming (OOP) and Functional Programming (FP) paradigms for privacy-preserving machine learning strategies (e.g., FedAvg, FedProx, and FedOpt). The source code for our method is publicly available at https://github.com/TaehwanY98/Fed-Ref.
@@ IV-A3 FeTs2022
To evaluate FedRef, we perform the segmentation task on the FeTs2022 dataset, which features a specific non-IID data partitioning, as shown in Fig. 3. The baseline model used for this task is a 3D U-Net.
IV-A4 CINIC-10
Table II compares the different types of client-side computations and the communicated resources for each method. FedRef clients perform local training and communicate model parameters and single cost values to the server. The combination of low client-side computational costs and low communication requirements enables efficient communication with fast local training.
@@ IV-A4 CINIC-10
TABLE III: FEMNIST classification: required rounds
Method $AsymmetricLoss<0.2$
FedAvg 21 rounds
FedProx 20 rounds
FedOpt 21 rounds
FedRef 20 rounds
TABLE IV: CINIC-10 classification: required rounds
Method $AsymmetricLoss<0.4$ $F1score>18\%$
FedAvg 28 rounds 22 rounds
FedProx 28 rounds 22 rounds
FedOpt 28 rounds 22 rounds
FedRef 23 rounds 23 rounds
TABLE V: FeTS2022 segmentation: required rounds
Method $mIOU>14\%$ $Hausdorff95<9$
FedAvg 30 rounds 23 rounds
FedProx 27 rounds 18 rounds
FedOpt 28 rounds 19 rounds
FedRef 27 rounds 17 rounds
In Tables III and IV, we applied a tiny convolutional neural network (tiny-CNN) as the baseline model for small-sized images to minimize overfitting on the FEMNIST and CINIC-10 datasets. As a result, our FedRef approach achieves a training speed comparable to other optimization strategies (FedProx, FedOpt) under the tiny-CNN setting, despite the reduced client-side computational costs.
@@
Furthermore, in Table V, FedRef achieves better model performance, with a higher mean intersection of union (mIoU) and a lower Hausdorff distance than other methods on medical images exhibiting data heterogeneity and data skew. Since the cost function corresponds to the Dice score relative to the evaluation metric (mIoU), FedRef effectively optimizes and minimizes the cost value. Consequently, the performance of the FedRef model is strongly influenced by the cost function.
@@
V Challenges
FedRef still faces several challenges, including the selection of hyperparameters (such as the client-server learning rate and the constant value representing the total number of selected models from previous rounds) and ensuring protection against malicious users and model inference attacks. As different assumptions can be made regarding the FIM, we plan to investigate various FIM formulations in future work to identify a more suitable proximal term.
VI Conclusions

### v4

@@ I Introduction
Critical issues and Bayesian optimization. In the existing studies section, we discuss FedAvg, FedProx, and FedOpt, which are existing approaches for model optimization in FL scenarios, along with their limitations regarding client-side computational costs and catastrophic forgetting [mccloskey1989catastrophic][french1999catastrophic] [goodfellow2014empirical]. We also explain Bayesian Transfer Learning. We then present FedRef, a novel, communication-efficient Bayesian fine-tuning approach that leverages a reference model. FedRef mitigates catastrophic forgetting in each round by integrating features from previous models while performing maximum a posteriori (MAP) estimation.
A definition of MAP estimation In earlier work, MAP estimation has been used to optimize an objective function for transfer learning between a target model with pre-trained parameters and a source model [bayesian]. In contrast, traditional model optimization in FL scenarios employs a proximal term to minimize the deviation of a local model from the global model [fedprox]. This proximal term can also be considered when performing MAP estimation. In our research, we use MAP estimation to optimize an objective function that is updated using aggregated model information from previous rounds exclusively on the server side, thereby minimizing the computational costs for clients.
Fig. 1 provides an overview of our proposed FedRef algorithm. The FedRef algorithm can be summarized in the following steps:
  1. 1.
    Server sends a model to clients.
  2. 2.
    Clients perform local training.
  3. 3.
    Clients send a model parameter and cost value.
  4. 4.
    Server aggregates selected clients parameters.
  5. 5.
    Server computes a reference model.
  6. 6.
    Server optimize global model derived from our Bayesian optimization using the reference model.
  7. 7.
    Server send the result of the optimization to the clients.
@@
For the segmentation task, we present the predictive model performance across different rounds using graphical representations, based on evaluation metrics commonly used in medical image segmentation studies [karimi2019reducing, chen2021transunet, hatamizadeh2022unetr, zhu2022medical]: mean Intersection over Union (mIoU), Dice Coefficient (DC), and Hausdorff Distance . And we present required rounds on reaching specific evaluation scores by Tables for each datasets and methods. And then finally, we demonstrate communication efficiency from our FedRef method by charts including relative communication resource ratios using required rounds on reaching specific evaluation scores.
In summary, the proposed FedRef:communication efficient Bayesian fine tuning with reference model method demonstrates high predictive performance on both medical and general computer vision tasks while minimizing computational costs for clients. Due to minimized computational costs and parameter-efficient fine-tuning (PEFT), We achieve and contribute higher communication efficiency on federated learning scenarios than existing studies.
Figure 1: Overview of the our FedRef: communication-efficient bayesian fine tuning using a reference model Algorithm.
II Existing studies
Federated Learning has emerged as a distributed machine learning paradigm enabling collaborative model training without centralized data collection, thus preserving user privacy. Meanwhile, parameter-efficient fine-tuning (PEFT) techniques allow efficient adaptation of large-scale pre-trained models using only a fraction of their parameters. The combination of FL and PEFT represents a promising direction for privacy-preserving and resource-effective learning across distributed devices.
For the federated learning backgrounds and paradigms, we introduce basic federated learning papers about optimizing model parameters and fine tuning.
@@ II-3 FedOpt
The authors of [reddi2020adaptive] focus on integrating adaptive optimizers such as Adam, Yogi, and Adagrad at the server side, leading to variants of FedAvg that are known as FedAdam, FedYogi, and FedAdagrad, respectively. Following the FedAdagrad equation (4):
@@
$g_{r}$: averaged update from all participating clients, $G_{r}$: accumulated sum of squared gradients (maintained on the server), $\tau$: small constant for numerical stability (e.g., $10^{-6}$).
Following the FedAdam Equation (II-3):
@@ II-4 Bayesian Transfer Learning
As a result, this paper investigates the optimization of transfer learning in an FL scenario. Whereas conventional transfer learning happens between only two models – one a pre-trained model intended for transfer, and the other a target mode – in FL scenarios, the server must aggregate and optimize many different client models. Therefore, we need a new MAP estimation problem formulation for FL scenarios. For completeness, the mathematical background on Laplacian regularization in Bayesian optimization and a comparativediscussion with FedDyn are provided in the Appendix.
@@ III FedRef: Communication-Efficient Bayesian Fine Tuning using a Reference Model
The existing studies have limitations of high client-side computational costs and catastrophic forgetting issues on each rounds. For optimal model performance and low client-side computational cost, we propose FedRef, a communication-efficient Bayesian fine-tuning method with a reference model that overcomes catastrophic forgetting by inferring from global models of previous rounds.
@@
III-A Descriptions of a reference model
As illustrated in Fig. 2, the reference model serves as a target model of proximal term objective reference to mitigate a deviation from previous round models. In the existing study [bayesian], transfer learning is a concept of model distillation and focus on how to distill and transfer features from alternative reference models to a target model. The most of FL scenario also focus on how do distillate to transfer features from each clients to global model. Therefore, in the most of FL scenarios, the ideal case assumes each client’s model inference is an independent event and when the model inference of all clients is, on average, the most accurate phase.
III-B A MAP estimation in FL scenarios
@@
$D_{ref}$ is defined solely to explain our optimal MAP problem in Equation (III-C). This approach helps overcome catastrophic forgetting in each round by integrating features from previous rounds into the MAP problem and subsequently optimizing the integrated MAP value.
III-C A objective estimation derived from the MAP estimation
Equation (11) reformulates global optimization in an FL setting as a MAP estimation, where the prior term $\log\,p(\theta|D_{ref})$ incorporates historical knowledge from previous global models, and the likelihood terms $\log\,p(D_{k}|\theta)$ accumulate evidence from the currently participating clients. Finally, the objective function derived in Equation (11) can be expressed by combining Equations (8, 10, and 12) to form Equation (III-C). Likelihood approximation via local loss is denoted at Equation 12,13. Equation (11) expresses the MAP objective in a federated setting, where the server combines a prior derived from a reference model with the likelihood contributions from all participating clients. Since the server cannot access raw client data, each client’s negative log-likelihood is approximated by its local loss function.
$-\log p(D_{k}\mid\theta)\approx F_{k}(\theta)$ (12)
Equation (12) formalizes this approximation, allowing the likelihood term to be replaced with a computable local loss.
$\sum_{k=1}^{K}\log p(D_{k}\mid\theta)\Longleftrightarrow-\sum_{k=1}^{K}F_{k}(\theta)$ (13)
Based on this, Equation (13) shows that the global likelihood can be written as a weighted sum of client losses, where the weights reflect the relative data sizes of each client. This produces a practical formulation that aligns with standard FL aggregation.
To incorporate knowledge from previous rounds, FedRef introduces a reference model that serves as the prior mean. Using a Laplace approximation, the prior becomes an $L_{2}$ regularization term encouraging the global model to remain close to this reference. This helps stabilize optimization and mitigates catastrophic forgetting in non-IID environments.
Finally, Equation (III-C) results from combining the likelihood approximation and the prior regularization. This objective balances client empirical loss with temporal consistency, enabling FedRef to perform stable and communication-efficient global updates.
@@ III-C A objective estimation derived from the MAP estimation
In Equation (III-C), $F_{k}$: local loss value from client $k$, the constant term $\sum_{k}^{K}F_{k}$ denotes the sum of client losses. The diagonal matrix $\mathrm{diag}(W_{1},\dots,W_{K})$ represents the aggregation weights (e.g., $\frac{n_{k}}{n}$). The $\theta^{1}$ is a global model and $\theta^{2}$ is a reference model. The term $(\theta_{r}-\theta_{r-1})^{2}$ signifies the $L_{2}$ regularization of the model parameters $\theta_{r}-\theta_{r-1}$.
III-D Formal definition of a reference model
@@
The resulting objective $L_{ref}$ thus combines the weighted sum of client-side losses with a temporally regularized penalty that constrains the drift of global parameters, effectively balancing adaptation and retention across communication rounds.
The server computes the gradient of $L_{ref}$ and performs a single Bayesian optimization step, as detailed in Algorithm 5 (lines 17–18). This step updates the global parameters $\theta^{1}$ by jointly considering aggregated client losses ($F_{k}$) and reference constraints ($\theta^{2}$).
@@ IV-A1 Environments
Within Python, the Flower framework enables the easy construction of federated learning servers and clients from high-level code. It supports both Object-Oriented Programming (OOP) and Functional Programming (FP) paradigms for privacy-preserving machine learning strategies (e.g., FedAvg, FedProx, and FedOpt). The source code for our method is publicly available at MyGithub.
@@ IV-A4 FeTs2022
To evaluate FedRef, we perform the segmentation task on the FeTs2022 dataset, which features a specific non-IID data partitioning, as shown in Fig. 3. The baseline model used for this task is a 3D U-Net [cciccek20163d]. 3D U‑Net provides a powerful framework for volumetric medical image segmentation, extending convolutional feature extraction into three dimensions to exploit full 3D context. Despite its computational demands, it has become a foundation for modern medical imaging pipelines and continues to inspire improved architectures for efficient and precise 3D segmentation.
IV-B Experimental Results
Evaluations of predictive model performances. Table II compares the different types of client-side computations and the communicated resources and server-side computations for each method. FedRef clients perform local training and communicate model parameters and cost values to the server. The combination of low client-side computational costs and low communication requirements enables efficient communication with fast local training.
@@ IV-B Experimental Results
An ablation study of FedRef. We conduct an ablation study on the two main hyperparameters of FedRef: the regularization strength $\lambda$ and the number of reference models $p$. As shown in Table III, $\lambda$ controls the influence of the reference model during optimization. Larger $\lambda$ values strengthen temporal regularization, and $\lambda$ between 0.0001 and 0.001 provides stable convergence across all datasets. For FEMNIST and CINIC‑10, $\lambda$ = 0.001 yields the best performance, while smaller $\lambda$ values perform slightly better on FeTS2022 due to stronger non‑IID conditions.
The parameter $p$ determines how many past global models are averaged to build the reference model. While larger $p$ increases temporal stability, it also raises memory usage. In practice, $p$ = 3 offers a good balance between performance and computational cost, and higher values do not yield additional improvements.
These results confirm that both $\lambda$ and $p$ play complementary roles: $\lambda$ governs the strength of reference-based regularization, and $p$ controls the temporal depth of the reference model. Proper tuning of these parameters enables FedRef to maintain stable optimization while preserving communication and computational efficiency.
TABLE III: FedRef heuristic parameters analysis: total round 30
$\lambda$ $p$ FEMNIST: AsymmLoss CINIC-10: AsymmLoss FeTs2022: DC FeTs2022: HF95
0.0001 3 13.95 9.245 14.09% 8.685
0.0001 5
0.0005 3 13.95 8.860 14.09% 8.676
0.0005 5
0.001 3 14.09 8.392 13.98% 8.847
0.001 5
Evaluations of communication efficiency. This section evaluates how efficiently FedRef reduces the communication cost required to reach a given performance level, compared to baseline optimization strategies such as FedAvg, FedProx, and FedOpt. Because FedRef does not require clients to compute or transmit proximal terms, the amount of data communicated per round remains similar to FedAvg. However, by performing Bayesian optimization with a reference model entirely on the server side, FedRef accelerates convergence and therefore lowers the total number of communication rounds needed. Fig. 4. reports the relative communication resource ratio, computed using Equation (17), for the FEMNIST, CINIC-10, and FeTS2022 benchmarks. This ratio incorporates both the number of bytes communicated per round ($b$) and the required rounds ($\gamma$) to reach predefined evaluation thresholds. A lower value indicates higher communication efficiency.
As summarized in Tables IVVI, FedRef consistently requires fewer rounds to reach the target performance across all benchmarks. On FEMNIST, FedRef requires 12 rounds to reach an Asymmetric Loss <14.5<14.5, outperforming FedAvg (13 rounds) and FedProx (16 rounds). On CINIC-10, FedRef achieves the Asymmetric Loss <9.5<9.5 threshold in 17 rounds, the best among all methods (1–3 rounds faster than others). On FeTS2022, FedRef reaches a Dice-Coefficient $14\%$ in 24 rounds, outperforming FedAvg (27 rounds) and achieving the target where FedOpt and FedProx fail.
TABLE IV: FEMNIST classification: required rounds
Method $AsymmetricLoss<14.5$ $Accuracy>9\%$
FedAvg 13 rounds 12 rounds
FedProx 16 rounds no meet the expectation
FedOpt 11 rounds 12 rounds
FedRef 12 rounds 12 rounds
TABLE V: CINIC-10 classification: required rounds
Method $AsymmetricLoss<9.5$ $Accuracy>10\%$
FedAvg 20 rounds 22 rounds
FedProx 19 rounds 22 rounds
FedOpt 18 rounds 22 rounds
FedRef 17 rounds 22 rounds
TABLE VI: FeTS2022 segmentation: required rounds
Method Dice-Coefficient $14\%$ Hausdorff95 <9<9
FedAvg 27 rounds 19 rounds
FedProx no meet the expectation 28 rounds
FedOpt no meet the expectation 20 rounds
FedRef 24 rounds 16 rounds
An evaluation metric definition. In the Fig. 4, relative communication resource ratios on each datasets are computed derived from equation (17):
$E_{max}(b,\gamma,\epsilon)=\text{max}(2\times b\times\gamma)$
$E_{min}(b,\gamma,\epsilon)=\text{min}(2\times b\times\gamma)$
$E=\frac{E_{s}-E_{min}}{E_{max}-E_{min}}+\epsilon$ (17)
The $E$ is a list of relative communication resource ratios on each federated learning strategies. $b$ includes the bytes of communication resource parameters (model parameter,… etc) on each federated learning strategies. $\gamma$ is a required rounds of specific standard of evaluation. $\epsilon$ can be set as a constant value for visualization in the Fig. 4. As a result, our proposed FedRef achieves higher communication efficient on reaching required rounds than existing works derived from the Fig. 4.
For further detail, in the Fig. 4, on FEMNIST, required round $\gamma$ represents $\gamma\in[11,12,13,16]$ derived from the Table IV if we define the standard of evaluation score to asymmetric loss. And $p$ includes 8.2 mega bytes from the global model. For the $\epsilon$ setting, we set $\epsilon$ to 1 in our experiments. In the Fig. 4, on CINIC-10 also, required round $\gamma$ represents $\gamma\in[17,18,19,20]$ derived from the Table V. And $p$ includes 8.3 mega bytes from the global model. For the $\epsilon$, we set $\epsilon$ to 1 in our experiments. In the Fig. 4, on FeTs2022 lastly, required round $\gamma$ represents $\gamma\in[16,19,20,28]$ derived from the Table V if we define the standard of evaluation score to mean Hausdorff distances. And $p$ includes 8.3 mega bytes from the global model. For the $\epsilon$, we set $\epsilon$ to 1 in our experiments.
@@
Summary of results. In Tables IV and V, we applied a tiny convolutional neural network (tiny-CNN) as the baseline model for small-sized images to minimize overfitting on the FEMNIST and CINIC-10 datasets. As a result, our FedRef approach achieves a training speed comparable to other optimization strategies (FedProx, FedOpt) under the tiny-CNN setting, despite the reduced client-side computational costs. Furthermore, in Table VI, FedRef achieves better model performance, with both higher mean intersection of union (mIoU) and mean dice-coefficient (DC) and a lower Hausdorff distance than other methods on medical images exhibiting data heterogeneity and data skew. Since the cost function corresponds to the Dice score relative to the evaluation metric (DC), FedRef effectively optimizes and minimizes the cost value. Therefore, the performance of the FedRef model is strongly influenced by the cost function.
Consequently, FedRef reduces the total communication burden without increasing client-side computation, offering a practical and scalable optimization strategy for real-world federated learning scenarios.
Figure 4: Relative communication resource ratios on each datasets
@@
V Discussion and Future Work
FedRef still faces several challenges, including the selection of hyperparameters (such as the client-server learning rate and the constant value representing the total number of selected models from previous rounds) and ensuring protection against malicious users and model inference attacks. As different assumptions can be made regarding the FIM, we plan to investigate various FIM formulations in future work to identify a more suitable proximal term. Moreover, for preventing unbounded drift issues from non-IID federated learning scenarios which can be described as a scenario where this divergence is not naturally limited or controlled, potentially leading to performance degradation, slow convergence, or even model collapse, to mitigate the issues, alternative regularization techniques can be a solutions of the unbounded drift issues. Moreover, while FedRef reduces the client-side burden, it implicitly assumes that the reference model can be reconstructed from a fixed number of previous rounds ($p$). In highly dynamic client participation scenarios, this assumption may not hold, potentially affecting the stability of the Bayesian objective. Future studies could therefore explore adaptive reference model selection or hierarchical priors to further stabilize optimization without sacrificing efficiency.
VI Related work
The parameter efficient fine tuning (PEFT) studies contributed to decrease required rounds achieving specific evaluation scores and to minimize various non-IID issues where is occurred from real world federated learning scenarios. In the existing PEFT on federated learning studies [fedprompt], model freezing and compression methods can increase communication efficiency. And alternative optimization studies [li2018hyperband][jaderberg2017pbt][chen2025fedshaplex] also can be references of alternative objective estimations against our a MAP estimation. Furthermore, the existing personalized federated learning studies [scott2023pefll][zhang2025fedlth] also can be references for better predictive model performance on each client cases. In terms of communication efficient federated learning, the existing studies [sezgin2025energy][shi2025federated] not only consider minimizing computational costs and communication resources and compressing model, but also minimizing client and server side energy consumptions. For the studies [dou2025toward] [zhao2021sear] of protecting the global model from malicious users and clients, can be references of solution about filtering malicious users and clients and regularization of the malicious users and clients.
VII Conclusions
@@
Acknowledgment
This research was supported by the MSIT, Korea, NRF Korea (RS-2025-00557379, 50%) and the Innovative Human Resource Development for Local Intellectualization support program (IITP-2026-RS-2022-00156360, 30%) supervised by the IITP, and Convergence Security Core Talent Training Business Support Program (IITP-2026-RS-2024-00426853, 20%) supervised by the IITP.
Appendix A. Laplacian Expression in Bayesian Optimization
The Laplacian Expression in Bayesian Optimization: Bayesian Optimization (BO) seeks the global optimum of an unknown and expensive-to-evaluate objective function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ by constructing a probabilistic surrogate model and iteratively selecting evaluation points to balance exploration and exploitation. A common surrogate is the Gaussian Process (GP), which assumes that function values follow a joint Gaussian distribution characterized by a mean function $m(\mathbf{x})$ and a covariance (kernel) function $k(\mathbf{x},\mathbf{x}^{\prime})$.
Laplacian-Regularized Acquisition Optimization: To encourage spatially smooth exploration, the acquisition function $\alpha(\mathbf{x})$ can be regularized by the Laplacian to penalize irregular updates:
$\mathcal{L}(\alpha)=\int_{\Omega}\left\|\nabla\alpha(\mathbf{x})\right\|^{2}\,d\mathbf{x}=-\int_{\Omega}\alpha(\mathbf{x})\,\Delta\alpha(\mathbf{x})\,d\mathbf{x},$ (18)
where $\Omega$ denotes the search region. Minimizing $\mathcal{L}(\alpha)$ enforces smooth acquisition surfaces and stabilizes optimization over noisy, high–dimensional domains.
Laplace Operator in Bayesian Linear and Partial Differential Equation(PDE)-Constrained Models: In physics-informed or PDE-constrained Bayesian optimization, the Laplacian explicitly appears in the governing equation:
$\mathcal{A}(f)(\mathbf{x})=\Delta f(\mathbf{x})-g(\mathbf{x})=0,$ (19)
where $\mathcal{A}$ is the differential operator defining the system. Here, BO aims to infer or optimize $f$ under this partial differential constraint. The corresponding GP prior can be constructed by designing a covariance function $k$ that satisfies the same PDE constraint, i.e.,
$\mathcal{A}_{\mathbf{x}}\,\mathcal{A}_{\mathbf{x}^{\prime}}\,k(\mathbf{x},\mathbf{x}^{\prime})=\delta(\mathbf{x},\mathbf{x}^{\prime}),$ (20)
where $\delta(\cdot,\cdot)$ is the Dirac delta function.
Laplacian-based Transfer Regularization: To preserve smoothness and avoid negative transfer, some transfer learning and BO formulations impose a Laplacian regularization term.
$\Omega(f)=\frac{1}{2}\sum_{i,j}w_{ij}\bigl(f^{(i)}-f^{(j)}\bigr)^{2}=\mathbf{f}^{\top}L\mathbf{f},$ (21)
where L=DWL=D-W is the graph Laplacian derived from the inter-task similarity matrix $W$. This penalty enforces function consistency across related tasks and integrates naturally with GP priors or acquisition optimization.
Appendix B. FedDyn: Federated Learning with Dynamic Regularization
FedDyn: Federated Learning with Dynamic Regularization: (Federated Dynamics) is an optimization framework proposed to mitigate the challenges of data heterogeneity in Federated Learning (FL). In conventional FL approaches such as Federated Averaging (FedAvg) [fedavg], model divergence arises when local client updates differ significantly due to non-i.i.d. data distributions. FedDyn addresses this issue by introducing a dynamic regularization term that implicitly aligns local and global objectives, ensuring stable and fast convergence under client heterogeneity.
FedDyn modifies the local objective of each client to introduce a dynamic correction term, yielding the following local problem:
$\min_{\mathbf{w}_{k}}\bigl(f_{k}(\mathbf{w}_{k})-\boldsymbol{\lambda}_{k}^{\top}\mathbf{w}_{k}+\tfrac{\alpha}{2}\|\mathbf{w}_{k}\|^{2}\bigr),$ (22)
where $\alpha$ is a tunable regularization coefficient and $\boldsymbol{\lambda}_{k}$ is a dynamically updated gradient correction variable.
The global update is given by:
$\mathbf{w}^{t+1}=\frac{1}{K}\sum_{k=1}^{K}\bigl(\mathbf{w}_{k}^{t+1}-\tfrac{1}{\alpha}\boldsymbol{\lambda}_{k}^{\,t}\bigr),$ (23)
followed by
$\boldsymbol{\lambda}_{k}^{\,t+1}=\boldsymbol{\lambda}_{k}^{\,t}-\alpha(\mathbf{w}^{t+1}-\mathbf{w}_{k}^{t+1}).$ (24)
This coupling of local and global variables introduces a potential game structure that ensures consistency between global and local optimization dynamics.
The FedDyn formulation can be understood as constructing a penalized Lagrangian system between the global and local models:
$L(\{\mathbf{w}_{k}\},\mathbf{w},\{\boldsymbol{\lambda}_{k}\})=\sum_{k}p_{k}f_{k}(\mathbf{w}_{k})+\frac{\alpha}{2}\sum_{k}p_{k}\|\mathbf{w}_{k}-\mathbf{w}\|^{2}.$ (25)
The dynamic variable $\boldsymbol{\lambda}_{k}$ accumulates discrepancies between local and global models, thereby compensating for non‑i.i.d. data effects. When aggregated, these correction terms implicitly align all local objectives with the global one, leading to faster and more robust convergence than standard FedAvg or FedProx.